# Metacognitive particles, mental action and the sense of agency

### v1

@@ 3.2 Cognitive and metacognitive particles
Note that in the active case, $\mu^{(1)}$ can constitute the entire set of parameters of the particle’s belief about the world. Whereas passive metacognitive beliefs are necessarily abouta subset of the parameters of the particle’s belief about the world.
@@ 7 Higher forms of metacognition
One can go further and consider multiply nested particles, that is a particle within a particle within a particle, etc. See Figure 4. Proceeding exactly as above, the internal paths at the highest level can encode beliefs about the internal paths of the level below and so forth. And the internal paths at the lowest level encode beliefs about external paths. In this case we have a particle with as many levels of metacognition as there are nested particles. At each level, the active path can influence the internal path at the level below, which itself influences the sensory path at the current level, which in turn influences the internal path, closing the (mental) action (mental) perception loop. The nested specification is equivalent to saying that the generative model $P(x)$ describing the system is a hierarchical model with as many levels as there are nested components. Finally, multiply nested particles may be $N$-times comprehensively metacognitive in the sense that at each level the internal paths hold beliefs about all of the internal paths at the level below (See Section 8.1 Equation 22).
@@ 8.3 Computational phenomenology
Perhaps more interestingly is the link between the theoretical treatment presented here and computational models of introspection and meta-awareness describing metacognitive processes such as meditation. As we attempt a mathematical definition of metacognition, this raises the question whether the description presented in this paper could be refined to get closer to our computational and empirical understanding of this phenomenon. Recent advances in active inference modelling of meditation suggest that some aspects can be adequately modelled as active inference under a hierarchical generative model where actions at the higher level modulate the precisions (i.e. inverse covariances) of beliefs encoded by the level below [sandved-smithComputationalPhenomenologyMental2021]. From this perspective, this amounts to the behaviour of multiply nested strange particles [fristonPathIntegralsParticular2023], where inner particles hold beliefs about active paths that influence the [sufficient statistics of the] beliefs held by the lower-level particle. There is an interesting road ahead where computational models of meditation furnish examples of what we would expect from metacognitive phenomenology, and theoretical treatments along these lines showing general conditions for the phenomenology described by these models to arise.
For example, a central theme in contemplative practices such Mahamudra or self-inquiry is the first person investigation of the apparent sense separation from the external world. In these practices, attention is directed towards the impression of perceiving things ‘over there’ from a perspective that is ‘in here’. In a similar way that the sense of agency can be related to the joint probability distribution over internal and active paths (Section 6), this sense of separation finds a potential mathematical candidate in the joint probability distribution over internal and external paths. This belief will necessarily encode the understanding that what is internal to me is separated from what is external, since the Markov blanket enforces a conditional independence of internal and external paths.
@@ 8.4 Empirical testing
It is possible to make empirical predictions in regards to the formalisation of the sense of agency. For example, we predict that the KL divergence in (20) will be zero, or low, in the absence of intentional binding [haggardSenseAgencyHuman2017].
Another possible vector of empirical investigation is to examine the nature of the influence of higher level dynamics on lower level paths via (mental) actions mediated by inner active paths. Notice that the mechanism of action between levels, of an active metacognitive particle, is to exert an influence on parameters of lower level beliefs. This has implications for the type of physical interactions that might enable such an influence. For example, in a metacognitive particle we would expect that the dynamics of higher level internal paths be at the origin of processes that impact the dynamics of lower level paths. In neurobiology this leads to predictions, for example, about the existence of neuronal populations whose activity is at the origin of neuromodulatory dynamics. In psychology, this might be read as establishing an attentional set or, more simply, attention [ainleyBodilyPrecisionPredictive2016, duquetteIncreasingOurInsular2017, limanowskiSeeingDarkGrounding2018, sandved-smithComputationalPhenomenologyMental2021, senguptaNeuronalGaugeTheory2016, veissiereThinkingOtherMinds2020].

### v2

@@ 3.2 Cognitive and metacognitive particles
Note that in the active case, $\mu^{(1)}$ can constitute the entire set of parameters of the particle’s belief about the world. In contrast, passive metacognitive beliefs are necessarily about a subset of the parameters of the particle’s belief about the world.
@@ 7 Higher forms of metacognition
One can go further and consider multiply nested particles, that is, a particle within a particle within a particle, etc. See Figure 4. Proceeding exactly as above, the internal paths at the highest level can encode beliefs about the internal paths of the level below and so forth. And the internal paths at the lowest level encode beliefs about external paths. In this case we have a particle with as many levels of metacognition as there are nested particles. At each level, the active path can influence the internal path at the level below, which itself influences the sensory path at the current level, which in turn influences the internal path, closing the (mental) action (mental) perception loop. The nested specification is equivalent to saying that the generative model $P(x)$ describing the system is a hierarchical model with as many levels as there are nested components. Finally, multiple nested particles may be $N$-times comprehensively metacognitive in the sense that at each level the internal paths hold beliefs about all of the internal paths at the level below (See Section 8.1 Equation 22).
@@ 8.3 Computational phenomenology
There is an interesting link between the theoretical treatment presented here and computational models of introspection and meta-awareness describing metacognitive processes such as meditation. As we attempt a mathematical definition of metacognition, this raises the question whether the description presented in this paper could be refined to get closer to our computational and empirical understanding of this phenomenon. Recent advances in active inference modelling of meditation suggest that some aspects can be adequately modelled as active inference under a hierarchical generative model where actions at the higher level modulate the precisions (i.e. inverse covariances) of beliefs encoded by the level below [sandved-smithComputationalPhenomenologyMental2021]. From this perspective, this amounts to the behaviour of multiply nested strange particles [fristonPathIntegralsParticular2023], where inner particles hold beliefs about active paths that influence the (sufficient statistics of the) beliefs held by the lower-level particle. There is an interesting road ahead where computational models of meditation furnish examples of what we would expect from metacognitive phenomenology, and theoretical treatments along these lines showing general conditions for the phenomenology described by these models to arise.
For example, a central theme in contemplative practices such Mahamudra or self-inquiry is the first person investigation of the apparent sense of separation from the external world. In these practices, attention is directed towards the impression of perceiving things ‘over there’ from a perspective that is ‘in here’. Similarly to how the sense of agency can be related to the joint probability distribution over internal and active paths (Section 6), this sense of separation finds a potential mathematical candidate in the joint probability distribution over internal and external paths. This belief will necessarily encode the understanding that what is internal to me is separated from what is external, since the Markov blanket enforces a conditional independence of internal and external paths.
@@ 8.4 Empirical testing
It is possible to make empirical predictions in regards to the formalisation of the sense of agency. For example, we predict that the KL divergence in (20) will correlate with the behaviour, guided by the sense of agency, in the squares task, as participants dynamically close the action-perception loop to determine which square they control [perrykkadEffectUncertaintyPrediction2021]. Interestingly, this may open avenues of research for differences in mental and developmental disorders, such as autism spectrum condition, where metacognition may be implicated.
Another possible avenue of empirical investigation is to examine the nature of the influence of higher level dynamics on lower level paths via (mental) actions mediated by inner active paths. Notice that the mechanism of action between levels, of an active metacognitive particle, is to exert an influence on parameters of lower level beliefs. This has implications for the type of physical interactions that might enable such an influence. For example, in a metacognitive particle we would expect that the dynamics of higher level internal paths be at the origin of processes that impact the dynamics of lower level paths. In neurobiology this leads to predictions, for example, about the existence of neuronal populations whose activity is at the origin of neuromodulatory dynamics. In psychology, this might be read as establishing an attentional set or, more simply, attention [ainleyBodilyPrecisionPredictive2016, duquetteIncreasingOurInsular2017, limanowskiSeeingDarkGrounding2018, sandved-smithComputationalPhenomenologyMental2021, senguptaNeuronalGaugeTheory2016, veissiereThinkingOtherMinds2020].